What is the area of a circle with circumference = 10 meters?
1
Expert's answer
2020-12-10T11:54:41-0500

We need to find radius of circle in order to compute area of circle.

We have given it`s length or:

length=10(m)="2\times\pi\times R", it is easy to find R from previous equation:

"R=\frac{10}{2\times\pi}" then we can find area through "area=\pi\times R^2", So:

"area=\pi\times R^2"="\frac{\pi\times 10^2}{4\times \pi^2}=7.95 m^2" , this is the area of given circle.